# Two-Time Relativistic Bohmian Model of Quantum Mechanics

### v8

@@ II.1 Some fundamental properties
    basically connected to the term $\frac{\partial^{2}R}{\partial t^{2}}-c^{2}\nabla{}^{2}R$, which never is zero in all space (for photons, we will find that it vanishes just in direction of motion). Moreover, the second addend inside the root never is small compared to 1, even in the case of non-relativistic approximation. For a free particle, for example, we will find that $V_{Q}$ oscillates with an amplitude $\frac{1}{2}m\gamma_{c}c^{2}$. This energy is high but it only has the effect of spreading the particle in a volume of space around states that are independent of the Quantum Potential. Such states - orbits and stationary solutions in a box are examples of them - can therefore be found by the standard treatment, ignoring the Quantum Potential.
@@ III Free particle
Note that this motion arises by self-interaction with its own wavefunction, without the need to evoke the antiparticle. By imposing $v_{i_{s_{MAX}}}=c$, we obtain $\Delta{\mathsf{S}}=\frac{\lambda}{f\sqrt{2(f+2)}}$, which generalizes the standard Zitterbewegung if f is of the order of unity. According to point 1 of the theory, the amplitude $\Delta{\mathsf{S}}$ represents then the minimum positional uncertainty in $s$ of the particle: only by means of a wave packet description is it possible to take into account larger positional uncertainties and smaller momentum uncertainties. The uncertainty principle is therefore
@@
$\frac{1}{2}m\gamma_{c}c^{2}\times\frac{\pi\lambda}{f\sqrt{2(f+2)\,}c}=\frac{\pi\hbar}{2\,f\sqrt{2(f+2)}\gamma_{c_{s}}}\sim\frac{\hbar}{\gamma_{c_{s}}}.$ (22)
@@ VIII TTBM and the problem of Time in Quantum Mechanics
Finally, we emphasize that the intrinsic time defined in TTBM leads directly to an unambiguous TEUR, given by eq. (22).

### v9

@@ II.1 Some fundamental properties
    basically connected to the term $\frac{\partial^{2}R}{\partial t^{2}}-c^{2}\nabla{}^{2}R$, which never is zero in all space (for photons, we will find that it vanishes just in direction of motion). Moreover, the second addend inside the root never is small compared to 1, even in the case of non-relativistic approximation. For a free particle, for example, we will find that $V_{Q}$ oscillates with an amplitude $\frac{1}{2}m\gamma_{c}c^{2}$. This energy is really high but it only has the effect of spreading the particle in a volume of space around states that are independent of the Quantum Potential. Such states - atomic orbits and stationary solutions in a box are examples of them - can therefore be found by the standard treatment, ignoring the Quantum Potential.
@@ III Free particle
By imposing $v_{i_{s_{MAX}}}=c$, we obtain $\Delta{\mathsf{S}}=\frac{\lambda}{f\sqrt{2(f+2)}}$, which generalizes the standard Zitterbewegung if f is of the order of unity; however, here this motion arises by self-interaction with its own wavefunction, without the need to evoke the antiparticle. According to point 1 of the theory, the amplitude $\Delta{\mathsf{S}}$ represents the minimum positional uncertainty in $s$ of the particle: only by means of a wave packet description is it possible to take into account larger positional uncertainties and smaller momentum uncertainties. The uncertainty principle is therefore
@@
$\frac{1}{2}m\gamma_{c}c^{2}\times\frac{\pi\lambda}{f\sqrt{2(f+2)\,}c}=\frac{\pi\hbar}{2\,f\sqrt{2(f+2)}\gamma_{c_{s}}}\sim\frac{\hbar}{2\gamma_{c_{s}}}.$ (22)
@@ VIII TTBM and the problem of Time in Quantum Mechanics
Finally, we emphasize that the intrinsic time defined in TTBM leads without ambiguity to a directional TEUR, given by (cf. eq. (22))
$\Delta t\ \Delta E\gtrsim\frac{\hbar}{2\gamma_{c_{s}}},$ (38)
where $\gamma_{c_{s}}$ is the Lorentz factor relative to the component of the classical velocity in the generic direction $s$.